# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ 5.1. Probabilistic Approaches
Most probabilistic approaches, including the Hidden Factors as Topics (HFT) (McAuley and Leskovec, 2013) and Rating-Boosted Latent Topics (RBLT) (Tan et al., 2016), share a common methodology that integrates the latent dimension in ratings with topics extracted from reviews. This hybrid approach is particularly effective in addressing data sparsity by enhancing the system’s ability to make accurate predictions even with limited user history. Both models leverage textual reviews to refine the recommendation process, achieving improved effectiveness across diverse datasets such as Amazon and Yelp.
One of the unique features of HFT (McAuley and Leskovec, 2013) is its ability of extracting the topics from the review that can explain the variation present in ratings and reviews. In addition, this model demonstrates rating prediction accuracy, particularly for new products and users with limited rating history, compared to matrix factorization and LDA-based models. It can also identify the most useful and informative reviews, ensuring interpretability. However, it faces scalability challenges due to its computational demands. In contrast, RBLT (Tan et al., 2016) bridges the advantage of the latent factor model and topic model to integrate the ratings and textual reviews, enhancing the recommendation accuracy. Similar to HFT, RBLT alleviates data sparsity and improves interpretability by exploring item recommendability and user preference distributions in a shared topic space.
@@
Similarly, the Aspect-Aware Latent Factor Model (ALFM) (Cheng et al., 2018b) correlates latent topics extracted from reviews with latent factors from ratings, focusing on identical aspects. It surpasses traditional models by avoiding direct one-to-one correspondence between topics and factors, which improves both accuracy and interpretability but depends heavily on the quality of the review content. The Matching Distribution by Reviews (MDR) (Shalom et al., 2019) model effectively combines deep learning with collaborative filtering for text processing, capturing the distribution of user-item matches found in reviews. It provides better prediction accuracy and reduced runtime across multiple datasets. Finally, the Topic Initialized Latent Factor Model (TIM) (Peña et al., 2020) enhances recommendations accuracy by initializing user and item embeddings within a topic space derived from reviews. This method not only speeds up model convergence but also improves interpretability by aligning embeddings with identifiable topics, offering a more transparent and understandable recommender system.
@@
While all these models demonstrate enhanced recommendation accuracy, their effectiveness varies by application domain. For instance, HFT and RBLT are particularly beneficial for new products and users in e-commerce, where user history might be limited. In contrast, NGMM and the ALFM show broader applicability by also enhancing recommendations in specialized markets and for unique item groups. The detailed summary of the probabilistic models is presented in Table 2.
@@
Table 2. The detailed summary of the Probabilistic Approaches
Category Models Year Architecture Features Datasets Evaluation Tasks
Probabilistic HFT(McAuley and Leskovec, 2013) 2013 LDA ratings & reviews amazon, yelp & beer MSE & MAE accuracy & interpretability
RBLT (Tan et al., 2016) 2016 LDA ratings & reviews amazon MSE, P & R accuracy & interpretability
ALFM(Cheng et al., 2018b) 2018 topic reviews amazon & yelp RMSE accuracy
NGMM(Deng et al., 2018) 2018 NN, GMM reviews amazon MSE feature extraction
MDR (Shalom et al., 2019) 2019 LDA reviews amazon & yelp MSE accuracy & interpretability
TIM (Peña et al., 2020) 2020 LDA, MF ratings & reviews amazon & tripAdvisor Recall accuracy & interpretability
5.2. Deep Learning-based Approaches
Deep learning, a subset of machine learning, involves training large neural networks to recognize patterns in data. Deep learning architectures such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, GRU (Gated Recurrent Unit), Attention Mechanism, GNN (Graph Neural Networks), and Autoencoders have revolutionized various fields, including image and speech recognition, natural language processing, and recently, recommender systems. In recommender systems, deep learning-based models have shown an exceptional ability to extract deep, non-linear relationships between users and items, significantly improving recommendation quality. These models can handle vast amounts of data, learning from user interactions, item attributes, and additional context to provide personalized recommendations.
@@ 5.2. Deep Learning-based Approaches
Review-based recommender systems, in particular, have benefited from deep learning models. By analyzing the rich textual data in reviews, these systems uncover nuanced user preferences and item features that traditional methods might miss. In this subsection, we briefly summarize state-of-the-art deep learning-based models and analyze the key features along with their effectiveness. Some papers may apply multiple deep learning techniques in their recommender systems, and in this case, we categorize them based on the main models they have used in their systems.
@@
Convolutional Neural Network-based Approaches: Convolutional Neural Network (CNN) based models typically excel in addressing challenges associated with sparse user-to-item interactions by extracting the contextual features from textual reviews. CNNs can identify and capture local features, such as phrases and sentences, that carrys significant semantic meaning. This capability allows these models to make more accurate predictions, even with limited data. Most CNN-based approaches, such as Convolutional Matrix Factorization (ConvMF) (Kim et al., 2016) and Deep Cooperative Neural Networks (DeepCoNN) (Zheng et al., 2017), combine traditional Matrix Factorization methods with CNN’s power to enhance the overall effectiveness of recommender systems.
@@
The Convolutional Matrix Factorization (ConvMF) (Kim et al., 2016) integrates CNNs with Probabilistic Matrix Factorization to effectively capture detailed context within item descriptions, moving beyond the limitation of traditional bag-of-word techniques. ConvMF is particularly noted for its ability to improve recommendation accuracy in conditions of extreme data sparsity. However, its complexity and the detailed benefits of using pre-trained word embeddings indicate potential areas for further refinement and optimization. Utilizing dual neural networks, the Deep Cooperative Neural Networks (DeepCoNN) (Zheng et al., 2017) learns item properties and user behaviors from review texts. Its shared layer mimics factorization machine techniques, allowing intricate interactions of latent factors between users and items. Expanding on the concept of network representation, the Translation-Based Network Representation Learning (TransNet) (Catherine and Cohen, 2017) introduces vertex interactions in network representation learning as translation operations. This method allows TransNet to capture richer semantic information, enhancing the Social Relation Extraction task and providing deeper insights into vertex relationships.
@@
A novel sentiment-guided review-aware recommendation model (SentiRec) (Hyun et al., 2018) efficiently processes review data by transforming reviews into sentiment-encoded vectors. This significantly reduces computational demands and captures essential sentiment details, enabling tailored recommendations based on emotional content. Additionally, the Word-Driven and Context-Aware Review Modeling for Recommendation (WCN-MF) (Wang et al., 2018) model enhances text analysis by integrating Deep Latent Dirichlet Allocation with CNNs, allowing for a detailed extraction of contextual features from reviews, which traditional pooling methods might miss. This method improves the model’s contextual understanding and recommendation relevance.
@@
The Weighted Aspect-Based Opinion Mining Using Deep Learning for Recommender Systems (AODR) (Da’u et al., 2020) focuses on extracting weighted, aspect-based opinions, effectively identifying the relative importance users assign to different product aspects, thereby aligning recommendations more closely with user preferences. A hybrid Neural network to Combine Textual information and Rating for recommendation( NCTR ) (Liu et al., 2021a) integrates CNNs with collaborative filtering to address data sparsity, uncovering latent user-item interactions, and enriching the model’s understanding of complex preference patterns. Finally, Deep Rating and Review Neural Network for Item Recommendation (DRRNN) (Xi et al., 2021) leverages both ratings and reviews during training, ensures a comprehensive preservation of semantic information, which enhances personalization and offers a more holistic understanding of user and item dynamics.
@@
Neural Attention-based Approaches: Attention mechanisms within recommender systems significantly advance traditional algorithms by focusing on the most relevant parts of the data. This method effectively differentiates crucial information from noise, which enhances the interpretability and relevance of the recommendations. Broadly, these models excel in managing the detailed information present in user reviews, which significantly boosts the specificity and personalization of recommendations.
@@
The Attentive Aspect Modeling for Review-aware Recommendation (AARM) (Guan et al., 2019) specifically tackles the aspect-sparsity issue prevalent in many reviews. By learning from both consistent and varying aspect connections between users and products, AARM significantly enhances the depth of understanding in user-item relationships. Meanwhile, Cross-Modality Mutual Attention (NRCMA), (Luo et al., 2021) utilizes a two-tower structure to synchronize user and item representations through mutual attention, effectively enhancing the match between user preferences and item characteristics, though it faces challenges in very dynamic environments.
Lastly, the Enhanced review-based rating prediction (ERP) (Wu et al., 2021) incorporates a mechanism to evaluate the general influence of reviewers, allowing it to differentiate and tailor recommendations based on how much trust to place in each review. This capability ensures that the recommendations are not only personalized but are also aligned with the most credible and relevant user feedback, enhancing the accuracy and relevance of the suggestions.
Together, these attention-based models enhance recommender systems by improving how user data is interpreted and utilized. While all models enhance traditional methods by incorporating advanced attention mechanisms, they vary in applications and challenges. A3NCF focuses on aspect-based ratings, AIN adapts to contextual variations, AARM addresses review aspect sparsity, and NRCMA bridges user-item representation gaps.
Integrated CNN and Attention-based Approaches: This approach combines the strengths of Convolutional Neural Networks (CNNs) and attention mechanisms to enhance the accuracy of the recommender systems. By utilizing CNNs, this approach effectively extracts and learns deep, complex features from textual data, such as user reviews or item descriptions. The incorporation of attention mechanisms further refines the process by selectively focusing on the most relevant parts of the data, ensuring that the most informative aspects are emphasized. This integrated approach enables a more detailed understanding of user preferences and item characteristics, leading to more accurate and personalized recommendations.
@@
Context-Aware User-Item Representation Learning (CARL) (Wu et al., 2019a) and Dual Attention Mutual Learning (DAML) (Liu et al., 2019a) both excel in dynamically adapting recommendations based on user and item interactions. CARL leverages CNNs enhanced with an attention layer to derive insights from both review and interaction data, enriching the contextual understanding necessary for precise personalization. Demonstrating this, DAML integrates local and mutual attention mechanisms, which balance the representation of users and items to ensure accuracy and relevance in its outputs. Neural Recommendation with Personalized Attention (NRPA) (Liu et al., 2019b) personalizes recommendations by focusing on the informativeness of reviews for different users and items. The Aspect-aware Neural Review Rating (ARP) (Wu et al., 2019b) focuses on aspect-level opinions in reviews, providing a detailed understanding of user preferences and item qualities. By using a collaborative learning framework to train both review-level and aspect-level rating predictors, ARP improves rating prediction accuracy, particularly for items with sparse ratings.
@@
Recurrent Neural Networks (RNNs)-based Approaches: RNN and their variants, such as LSTM, and GRUs, are particularly effective in review-based recommender systems. By leveraging their ability to process sequential information, these models can analyze the detailed sentiment and progression within user reviews. This facilitates a deeper understanding of preferences and item characteristics from textual feedback. Consequently, RNNs enable recommender systems to generate more accurate and personalized recommendations by addressing complex details in user feedback.
The Sentiment-aware Deep Recommender System with neural Attention network (SDRA) (Da’u and Salim, 2019) further enhances the recommendations by fully leveraging user-generated textual reviews. Integrating a semi-supervised topic model with a deep learning framework (e.g. LSTM) through a neural attention mechanism, it adeptly captures both product aspects and the associated user sentiments from textual reviews, leading to improved recommendation performance. The use of Long Short-Term Memory (LSTM) encoders for capturing contextual word information and a co-attention mechanism for fine-tuning user-item interactions based on these insights further enhances performance. SDRA offers a pathway for incorporating user sentiment and aspect information into the recommendation process, showcasing its potential for more nuanced and effective recommendations. The Topical Attention Regularized Matrix Factorization (TARMF) (Lu et al., 2018) integrates attention-based GRUs with matrix factorization, tackling the inherent sparsity in user-item ratings that frequently challenges collaborative filtering systems. TARMF’s hybrid approach not only enhances sentiment analysis of product features but also adds a layer of transparency and explainability, setting it apart from traditional matrix factorization methods. The Selective Distillation Network (SDNet) (Chen et al., 2018b) introduces a distillation approach where the system uses external knowledge, such as user reviews, to train a “teacher” model that guides a more streamlined “student” model that focuses on rating prediction from user-item pair. The “teacher” model uses both CNN and LSTM for review processing to extract the enriched features from review. This separation allows for efficient runtime operation without sacrificing the depth of knowledge extraction, addressing the dual needs of performance and efficiency in dynamic environments.

### v3

@@ 5.1. Probabilistic and Topic Modeling Approaches
Most probabilistic approaches, such as the Hidden Factors as Topics (HFT) (McAuley and Leskovec, 2013) and Rating-Boosted Latent Topics (RBLT) (Tan et al., 2016), share a common methodology that integrates the latent dimension in ratings with topics. This hybrid approach is particularly effective in addressing data sparsity by enhancing the system’s ability to make accurate predictions even with limited user history. Both models leverage reviews to refine the recommendation process, achieving improved effectiveness across diverse datasets such as Amazon and Yelp.
One of the unique features of HFT (McAuley and Leskovec, 2013) is its ability to extract the topics from the review that can explain the variation present in ratings and reviews. In addition, this model demonstrates a high prediction accuracy, particularly for new products and users with limited rating history, compared to matrix factorization and LDA-based models. It can also identify the most useful and informative reviews, ensuring interpretability. However, it faces scalability challenges due to its computational demands. In contrast, RBLT (Tan et al., 2016) bridges the advantage of the latent factor model and topic model to integrate the ratings and textual reviews, enhancing the recommendation accuracy. Similar to HFT, RBLT alleviates data sparsity and improves interpretability by exploring item recommendability and user preference distributions in a shared topic space.
@@
The Aspect-Aware Latent Factor Model (ALFM) (Cheng et al., 2018b) correlates latent topics extracted from reviews with latent factors from ratings, focusing on identical aspects. It surpasses traditional models by avoiding direct one-to-one correspondence between topics and factors, which improves both accuracy and interpretability, but it depends heavily on the quality of the review content. The Matching Distribution by Reviews (MDR) model (Shalom et al., 2019) effectively combines deep learning with collaborative filtering for text processing, capturing the distribution of user-item matches found in reviews. It provides better prediction accuracy and reduced runtime across multiple datasets. Finally, the Topic Initialized Latent Factor Model (TIM) (Peña et al., 2020) enhances recommendations accuracy by initializing user and item embeddings within a topic space derived from reviews. This method not only speeds up model convergence but also improves interpretability by aligning embeddings with identifiable topics, offering a more transparent and understandable recommender system.
Table 2. The Detailed Summary of the Probabilistic Approaches
Models Year Methodology Features Datasets Evaluation Tasks
HFT (McAuley and Leskovec, 2013) 2013 LDA ratings & reviews amazon, yelp & beer MSE & MAE accuracy, interpretability
LMLF (Almahairi et al., 2015) 2015 LDA reviews amazon MSE accuracy
LBSN (Li et al., 2015) 2015 LDA reviews yelp & tripadvisor MSE accuracy
RBLT (Tan et al., 2016) 2016 LDA ratings & reviews amazon MSE, P & R accuracy
ParVecMF (Alexandridis et al., 2019) 2019 PMF ratings & reviews amazon, yelp MAP, MRR accuracy
ALFM (Cheng et al., 2018b) 2018 LDA reviews amazon & yelp RMSE accuracy, interpretability
NGMM (Deng et al., 2018) 2018 NN, GMM reviews amazon MSE accuracy
ARAOE (Hernández-Rubio et al., 2019) 2018 LDA reviews amazon, yelp P, Recall, nDCG accuracy, time complexity
MDR (Shalom et al., 2019) 2019 LDA reviews amazon & yelp MSE accuracy, interpretability
TIM (Peña et al., 2020) 2020 LDA, MF ratings & reviews amazon & tripadvisor Recall accuracy, interpretability
@@
While all these models demonstrate enhanced recommendation accuracy, their effectiveness varies by application domains. For instance, HFT and RBLT are particularly suitable for new products and users in the e-commerce domain, where user history might be limited. In contrast, NGMM and ALFM show broader applicability by also enhancing recommendations in specialized markets and for unique item groups. The detailed summary of the probabilistic models is presented in Table 2.
5.2. Deep Learning-based Approaches
Deep learning, a subset of machine learning, involves training large neural networks to recognize patterns in data. The idea of a neural network which is based on neurons came from biological neurons. A simple neural network is made up of multiple layers, and each layer has multiple neurons. The initial and final layers of a network are defined as input and output layers where layers in between are referred to as hidden layers. Neurons in each layer are connected to successive layers and they collect signals from previous layers. The signals on each neuron coming from the previous neurons are accumulated before transmitting to the next layer, and an activation function controls the output of each neuron (Goodfellow et al., 2016). The activation function processes the data in a non-linear fashion, making deep learning models capable of learning a complex representation of data. It is a widely used method of learning the representation of text, images, audio, and video signals. An architecture of a simple neural network also called Multi-Layer Perceptrons (MLPs) is shown in Figure 3.
[…]
A Simple Multi-Layer Perceptron (MLP) with 3 Layers
Figure 3. A Simple Multi-Layer Perceptron (MLP) with 3 Layers
Deep learning architectures such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory (LSTM) networks, Gated Recurrent Unit (GRU), Attention mechanisms, Graph Neural Networks (GNN), Large Language Models (LLMs) and Autoencoders have revolutionized various fields, including image and speech recognition, natural language processing, and recently, recommender systems. In recommender systems, deep learning-based models have shown an exceptional ability to extract deep, non-linear relationships between users and items, significantly improving recommendation quality. These models can handle vast amounts of data, learning from user interactions, and contextual information to provide personalized recommendations.
[…]
Summary of the Different Review-based Recommender Models.
@@ A Simple Multi-Layer Perceptron (MLP) with 3 Layers
By analyzing the rich textual data in reviews, these systems uncover nuanced user preferences and item features that traditional methods might miss. In this subsection, we briefly summarize state-of-the-art deep learning-based models and analyze the key features along with their effectiveness. Some papers may apply multiple deep learning techniques in their recommender systems, and in this case, we categorize them based on the main models they have used.
5.2.1. Convolutional Neural Network (CNN)-based Approaches:
@@
CNN-based models typically excel in addressing challenges associated with sparse user-to-item interactions by extracting the contextual features from textual reviews. CNNs can identify and capture local features, such as phrases and sentences, that carry significant semantic meaning. This capability allows these models to make more accurate predictions, even with limited data. Most CNN-based approaches, such as (Kim et al., 2016; Zheng et al., 2017), combine traditional Matrix Factorization methods with CNN’s power to enhance the overall effectiveness of recommender systems.
@@
The Convolutional Matrix Factorization (ConvMF) (Kim et al., 2016) integrates CNNs with probabilistic matrix factorization to effectively capture detailed context within item descriptions, moving beyond the limitation of traditional bag-of-word techniques. ConvMF is particularly noted for its ability to improve recommendation accuracy in conditions of extreme data sparsity. However, its complexity and the use of pre-trained word embeddings indicate potential areas for further refinement and optimization. Utilizing dual neural networks, the Deep Cooperative Neural Networks (DeepCoNN) (Zheng et al., 2017) learns item properties and user behaviors from review texts. Its shared layer mimics factorization machine techniques, allowing to capture the interactions between latent factors of users and items. Expanding on the concept of network representation, the Translation-Based Network Representation Learning (TransNet) (Catherine and Cohen, 2017) introduces vertex interactions in network representation learning as translation operations. This method allows TransNet to capture richer semantic information, enhancing the social relation extraction task and providing deeper insights into vertex relationships.
@@
A novel sentiment-guided review-aware recommendation model (SentiRec) (Hyun et al., 2018) efficiently processes review data by transforming reviews into sentiment-encoded vectors. This significantly reduces computational demands and captures essential sentiment details, enabling tailored recommendations based on emotional content. Additionally, the Word-Driven and Context-Aware Review Modeling for Recommendation (WCN-MF) model (Wang et al., 2018) enhances text analysis by integrating Deep Latent Dirichlet Allocation with CNNs, allowing for detailed extraction of contextual features from reviews, which traditional pooling methods might miss. This method improves the model’s contextual understanding and recommendation relevance.
@@
The Weighted Aspect-based Opinion Mining using Deep Learning for Recommender Systems (AODR) (Da’u et al., 2020) focuses on extracting weighted, aspect-based opinions, effectively identifying the relative importance users assign to different product aspects, thereby aligning recommendations more closely with user preferences. A hybrid Neural network to Combine Textual information and Rating for recommendation (NCTR) (Liu et al., 2021a) integrates CNNs with collaborative filtering to address data sparsity, uncovering latent user-item interactions, and enriching the model’s understanding of complex preference patterns. Finally, Deep Rating and Review Neural Network for Item Recommendation (DRRNN) (Xi et al., 2021) leverages both ratings and reviews during training, ensures a comprehensive preservation of semantic information, which enhances personalization.
@@
5.2.2. Neural Attention-based Approaches:
Attention mechanisms within recommender systems significantly advance traditional algorithms by focusing on the most relevant parts of the data. This method effectively differentiates crucial information from noise, which enhances the interpretability and relevance of the recommendations. Broadly, these models excel in managing the detailed information present in user reviews, which significantly boosts the specificity and personalization of recommendations.
@@ 5.2.2. Neural Attention-based Approaches:
The Attentive Aspect Modeling for Review-aware Recommendation (AARM) (Guan et al., 2019) specifically tackles the aspect-sparsity issue prevalent in many reviews. By learning from both consistent and varying aspect connections between users and products, AARM significantly enhances the depth of understanding in user-item relationships. Meanwhile, Cross-Modality Mutual Attention (NRCMA), (Luo et al., 2021) utilizes a two-tower structure to synchronize user and item representations through mutual attention, effectively enhancing the match between user preferences and item characteristics, though it faces challenges in very dynamic environments. The review-based recommender with attentive properties (RRAP) (Lei et al., 2023) integrates review properties and sequential information to improve recommendation quality. This model effectively reduces prediction error and enhances interpretability by leveraging attention mechanisms. However, the model relies on historical data and struggles with unseen reviews, which can limit its generalizability.
Lastly, the Enhanced review-based Rating Prediction (ERP) (Wu et al., 2021) incorporates a mechanism to evaluate the general influence of reviewers, allowing it to differentiate and tailor recommendations based on how much trust to place in each review. This capability ensures that the recommendations are not only personalized but are also aligned with the most credible and relevant user feedback, enhancing the accuracy and relevancy.
These attention-based models enhance recommender systems by improving how user data is interpreted and utilized. While all models enhance traditional methods by incorporating advanced attention mechanisms, they vary in applications and challenges. A3NCF focuses on aspect-based ratings, AIN adapts to contextual variations, AARM addresses review aspect sparsity, and NRCMA bridges user-item representation gaps.
5.2.3. Integrated CNN and Attention-based Approaches:
This approach combines the strengths of CNNs and attention mechanisms to enhance the accuracy of the recommender systems. By utilizing CNNs, this approach effectively extracts and learns deep, complex features from textual data, such as user reviews or item descriptions. The incorporation of attention mechanisms further refines the process by selectively focusing on the most relevant parts of the data, ensuring that the most informative aspects are emphasized. This integrated approach enables a more detailed understanding of user preferences and item characteristics, leading to more accurate and personalized recommendations.
@@ 5.2.3. Integrated CNN and Attention-based Approaches:
Context-Aware User-Item Representation Learning (CARL) (Wu et al., 2019a) and Dual Attention Mutual Learning (DAML) (Liu et al., 2019a) both excel in dynamically adapting recommendations based on user and item interactions. CARL leverages CNNs enhanced with an attention layer to derive insights from both review and interaction data, enriching the contextual understanding necessary for precise personalization. DAML integrates local and mutual attention mechanisms, which balance the representation of users and items to ensure accuracy and relevance in its outputs. Neural Recommendation with Personalized Attention (NRPA) (Liu et al., 2019b) personalizes recommendations by focusing on the informativeness of reviews for different users and items. The Aspect-aware Neural Review Rating (ARP) (Wu et al., 2019b) focuses on aspect-level opinions in reviews, providing a detailed understanding of user preferences and item qualities. By using a collaborative learning framework to train both review-level and aspect-level rating predictors, ARP improves rating prediction accuracy, particularly for items with sparse ratings.
@@
5.2.4. Recurrent Neural Networks (RNNs)-based Approaches:
RNN and their variants, such as LSTM, and GRUs, are particularly effective in review-based recommender systems. By leveraging their ability to process sequential information, these models can analyze the detailed sentiment and progression within user reviews. This facilitates a deeper understanding of preferences and item characteristics from textual feedback. Consequently, RNNs enable recommender systems to generate more accurate and personalized recommendations by addressing complex details in user feedback.
The Sentiment-aware Deep Recommender System with neural Attention network (SDRA) (Da’u and Salim, 2019b) further enhances the recommendations by fully leveraging user-generated textual reviews. Integrating a semi-supervised topic model with a deep learning framework (e.g., LSTM) through a neural attention mechanism, it adeptly captures both product aspects and the associated user sentiments from textual reviews, leading to improved recommendation performance. The use of LSTM encoders for capturing contextual word information and a co-attention mechanism for fine-tuning user-item interactions based on these insights further enhances performance. SDRA offers a pathway for incorporating user sentiment and aspect information into the recommendation process, showcasing its potential for more nuanced and effective recommendations. The Topical Attention Regularized Matrix Factorization (TARMF) (Lu et al., 2018) integrates attention-based GRUs with matrix factorization, tackling the inherent sparsity in user-item ratings that frequently challenges collaborative filtering systems. TARMF’s hybrid approach not only enhances sentiment analysis of product features but also adds a layer of transparency and explainability, setting it apart from traditional matrix factorization methods. The Selective Distillation Network (SDNet) (Chen et al., 2018b) introduces a distillation approach where the system uses external knowledge, such as user reviews, to train a “teacher” model that guides a more streamlined “student” model that focuses on rating prediction from user-item pair. The “teacher” model uses both CNN and LSTM for review processing to extract the enriched features from review. This separation allows for efficient runtime operation without sacrificing the depth of knowledge extraction, addressing the dual needs of performance and efficiency in dynamic environments.